In a basin area where oil is likely to be found underneath the surface, there are 3 locations with 3 different types of earth compositions, say C1, C2, & C3. The probabilities for these 3 compositions are 0.5, 0.3, & 0.2 respectively. Further, it has been found from the past experience that after drilling of well at these locations, the probabilities of finding oil is 0.2, 0.4, & 0.3 respectively. Suppose, a well is drilled at a location, & it yields oil, what is the probability that the earth composition was C1?



1
Expert's answer
2021-01-28T05:24:46-0500

Let "E_1,E_2,E_3" be three events such that , "E_1=" The location with earth composition was "C_1"

"E_2=" The location with earth composition was "C_2"

"E_3=" The location with earth composition was "C_3"

Also let "A" be an event such that , "A=" finding oil at the location

Then "P(E_1)=0.5,P(E_2)=0.3,P(E_3)=0.2" and "P(A\/E_1)=0.2,P(A\/E_2)=0.4,P(A\/E_3)=0.3"

Now we have to find "P(E_1\/A)."

According to Bayes' Theorem we know that , "P(E_1/A)=\frac{P(E_1).P(A/E_1)}{P(E_1).P(A/E_1)+P(E_2).P(A/E_2)+P(E_3).P(A/E_3)}"

"\therefore" The required probability is "P(E_1/A)=\frac{(0.5×0.2)}{(0.5×0.2)+(0.3×0.4)+(0.2×0.3)}=\frac{10}{28}=\frac{5}{14}"